Write a function that takes two numbers and perform all arithmetic operations.

Function should have default arguments set to 1, 1

● Addition

● Subtraction

● Division

● Multiplication


1
Expert's answer
2021-10-12T00:36:43-0400
def arithmetic(x=1, y=1):
    print('x+y =', x + y)
    print('x-y =', x - y)
    print('x*y =', x * y)
    print('x/y =', x / y)